If a probability of a win is 0.5 and the probability of a draw is 0.3, what is the probability of a loss?
The total has to add up to 1 in terms of probability, so the decimal has to be 0.2.
0.5 + 0.3 + 0.2 = 1
Answer: 0.2
Step-by-step explanation: Adding the probability of a win and a draw, you would get 0.8.
0.5+0.3=0.8
The probability of a loss is going to be the rest. Subtract 0.8 from 1 to get the probability.
1-0.8=0.2
The probability of a loss is 0.2.